Given a sequence an=((n!)2)/(2n)! . Evaluate lim n approaches infinity an

1
Expert's answer
2021-06-14T09:22:27-0400

"\text{By ratio test, }\\
If \space lim_{n→∞}\frac{a_{n+1}}{a_n}=L\space and\space L<1 \space then \space lim_{n→∞}a_n=0 .\\
Here, \\
a_n=\frac{(n!)^2}{(2n)!}\\
lim_{n→∞}\frac{a_{n+1}}{a_n}=lim_{n→∞}\frac{\frac{((n+1)!)^2}{(2n+2)!}}{\frac{(n!)^2}{(2n)!}}\\
=lim_{n→∞}\frac{((n+1)!)^2(2n!)
}{(2n+2)!(n!)^2}\\
=lim_{n→∞}\frac{((n+1).n!)^2(2n!)
}{(2n+2).(2n+1).(2n)!(n!)^2}\\
=lim_{n→∞}\frac{(n+1)^2
}{(2n+2).(2n+1)}\\
\text{Taking }n^2\text{ common from numerator and denominator.}\\
=lim_{n→∞}\frac{(1+\frac{1}{n})^2
}{(2+\frac{2}{n}).(2+\frac{1}{n})}\\
=\frac{1}{4}"